1)
The management of River Corporation is considering the purchase of a new machine costing $380,000 The companys desired rate of return is 6%. The present value factor for an annuity of $1 at interest of 6% for 5 years is 4.212. In addition to the foregoing information, use the following data in determining the acceptability of this investment: ncome from Net Cash Year Flow $20,000 20,000 20,000 20,000 20,000 $95,000 95,000 95,000 95,000 95,000 4 The average rate of return for this investment is Ja, 10.5% Ob. 25% oc. 15% od, 5%
2)
In capital rationing, alternative proposals that survive initial and secondary screening are normally evaluated in terms of: Oa. nonfinancial factors b. present value Oc. maximum cost d. net cash flow
3)
Below is a table for the present value of $1 at compound interest 6% 0.943 0.890 0.840 0.792 0.747 10% 0.909 0.826 0.751 0.683 0.621 12% 0.893 0.797 0.712 0.636 0.567 Year 2 4 Below is a table for the present value of an annuity of $1 at compound interest 6% 0.943 1.833 2.673 3.465 4.212 10% 0.909 1.736 2.487 3.170 3.791 12% 0.893 1.690 2.402 3.037 3.605 Year 4 Using the tables above, what would be the internal rate of return of an investment of $210,600 that would generate an annual cash inflow of $50,000 for the next 5 years? Од. 12% Ob. 14% С. 6% Od. 10%
4)
The process by which management allocates available investment funds among competing investment proposals is called Oa. investment capital Ob. investment rationing Oc. capital rationing Od.cost-volume-profit analysis
5)
The amount of the estimated average income for a proposed investment of $87,000 in a fixed asset, giving effect to depreciation (straight-line method), with a useful life of 4 years, no residual value, and an expected total income yield of $20,500, is a. $20,500 Ob. $5,125 Oc. $21,750 Od. $23,000
Answer #1
1 Option A is the Correct answer
Average rate of Return = Average Operating profit / Average Investment
Average Operating Profit = 20000
Average Investment = 380000/2
= 190000
Average Rate of Return = 20000 / 190000
= 10.5%
2 Option A is the Correct answer
In capital rationing, alternative proposals that survive initial and secondary screening are normally evaluated in terms of nonfinancial factors
3 Option C is the Correct answer
NPV at 6% for $ 1 = 4.212
NPV for $ 50000 for next 5 years @ 6% interest = 50000*4.212
210600
NPV of the transaction = NPV of 50000 for 5 years - Initial Investment
= 210600-210600
= 0
IRR is the rate where the NPV is 0
Internal Rate of Return 6%
4 Option C is the Correct answer
Capital Rationing is the process by which management allocates available investment funds among competing investment proposals.
5 Option B is the Correct answer
Expected Income 20500
Estimated Life 4
Average Investment for the period = 20500/4
Average Investment for the period 5125